OpenStudy (anonymous):

What is the maximum volume of a pyramid that can fit inside a cube that has side 18 cm long? A) 5832 cm3 B) 2916 cm3 C) 1944 cm3 D) 972 cm3

OpenStudy (asadkarim7):

18 ^3 = 5832 vol of cube

OpenStudy (saifoo.khan):

C is yr ans/

OpenStudy (asadkarim7):

B coz A cannot totally fit the cuce

OpenStudy (saifoo.khan):

the form is: 1/3 b^2 x h

OpenStudy (anonymous):

vol of pyramis = (1/3) * 18^2 * 18 =1944 C

OpenStudy (saifoo.khan):

Yay! i was Correct! :D

OpenStudy (anonymous):

lol Thanks guys!

OpenStudy (asadkarim7):

hey but its saying that what vol of pyramid that it could fit in

OpenStudy (asadkarim7):

the vol of cube would be 5832

OpenStudy (anonymous):

i think its a pyramid with height of 18 and base = area of base of the cube

OpenStudy (anonymous):

yep its definitely C

OpenStudy (saifoo.khan):

http://0.tqn.com/d/math/1/0/H/F/sbpyramidr.gif

OpenStudy (saifoo.khan):

asad^

OpenStudy (asadkarim7):

how d we know what shape is of base of pyramid

OpenStudy (saifoo.khan):

coz its a CUBE.

OpenStudy (anonymous):

it has to be - as asaifoo has said its inside a cube

OpenStudy (anonymous):

saifoo i meant

OpenStudy (saifoo.khan):

max vol = max base.

OpenStudy (saifoo.khan):

thnks math kid. (:

OpenStudy (anonymous):

Thanks to all of you for your eagerness to help!

OpenStudy (amistre64):

volume of a pyramid is: 1/3 (basearea) (height)

OpenStudy (amistre64):

base area here is 18^2; height is 18; and 1/3 is well ... 1/3

OpenStudy (amistre64):

im going with 1944

OpenStudy (saifoo.khan):

C C C

OpenStudy (saifoo.khan):

:D